OpenStudy (anonymous):

use the rack of 15 billiard balls shown. If one ball is selected at random, determine the odds in favor of it containing two digits.

OpenStudy (anonymous):

are the billiard balls numbered 1 - 15 ?

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

so there are 9 balls marked with 1 digit and 6 marked with 2 digit numbers.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Probability to select a ball at random, and it having a 2 digits is (number of balls with 2 digits) / (total number of balls) = 6/15=2/5=0.4